A research scholar wants to know how many times per hour a certain strand of virus reproduces. He wants to construct the 85%confidence interval with a maximum error of 0.19 reproductions per hour. Assuming that the mean is 12.3 reproductions and the standard deviation is known to be 2.5, what is the minimum sample size required for the estimate? Round your answer up to the next integer.
Answer #1

Given ME = 5, \sigma = 2.5, \alpha = 0.15

The Zcritical at \alpha = 0.05 is 1.44

The ME is given by :

critical_

Squaring both sides we get: (ME)2 = (Z critical)2 * \sigma2/n

Therefore n = (Zcritical * \sigma/ME)2 = (1.44*2.5/0.19)2 = 359.0027

Therefore n = 360 (Taking it to the next whole number)